# ChunkKV: Semantic-Preserving KV Cache Compression for Efficient Long-Context LLM Inference

### v1

@@ Abstract
To reduce memory costs in long-context inference with Large Language Models (LLMs), many recent works focus on compressing the key-value (KV) cache of different tokens. However, we identify that the previous KV cache compression methods measure token importance individually, neglecting the dependency between different tokens in the real-world language characterics. In light of this, we introduce ChunkKV, grouping the tokens in a chunk as a basic compressing unit, and retaining the most informative semantic chunks while discarding the less important ones. Furthermore, observing that ChunkKV exhibits higher similarity in the preserved indices across different layers, we propose layer-wise index reuse to further reduce computational overhead. We evaluated ChunkKV on cutting-edge long-context benchmarks including LongBench and Needle-In-A-HayStack, as well as the GSM8K and JailbreakV in-context learning benchmark. Our experiments with instruction tuning and multi-step reasoning (O1 and R1) LLMs, achieve up to 10% performance improvement under aggressive compression ratios compared to existing methods.
Large Language Models, KV Cache Compression, Model Evaluation
[…]
@@ 1 Introduction
Large Language Models (LLMs) have become essential for addressing various downstream tasks of natural language processing (NLP), including summarization and question answering, which require the interpretation of a long context from sources such as books, reports, and documents, often encompassing tens of thousands of tokens (Brown et al., 2020; Tay et al., 2022; Touvron et al., 2023). Recent advances in long-context technology within the field of machine learning (ML) systems (Dao, 2024; Jacobs et al., 2023; Xiao et al., 2024) have significantly enhanced computational throughputs and reduced latency of LLMs to process increasingly large input context lengths (Liu et al., 2024b; Young et al., 2024) with saving historical KV cache (key value attentions). However, the memory requirement of the KV cache in serving super-long contexts becomes a new bottlneck (Zhang et al., 2023; Reid et al., 2024). For instance, the KV cache for a single token in a 7B-parameter model requires approximately 0.5 MB of GPU memory, resulting in a 10,000-token prompt consuming around 5 GB of GPU memory.
@@
To address the substantial GPU memory consumption caused by KV caching, recent studies consider compressing the KV cache by pruning non-important discrete parts from the prompt tokens (Zhang et al., 2023; Li et al., 2024; Ge et al., 2023; Cai et al., 2024; Fu et al., 2024a; Yang et al., 2024b; Liu et al., 2024e; Tang et al., 2024). H2O (Zhang et al., 2023) and SnapKV (Li et al., 2024) have shown that retaining less than 50% of the discrete KV cache can significantly reduce GPU memory usage with minimal impact on performance. However, we identify that the previous KV cache compression methods (Zhang et al., 2023; Cai et al., 2024) measure token importance isolatedly, neglecting the dependency between different tokens in the real-world language characterics. For example, as shown in Figure 1, focusing on token-level importance might excessively focus on words about subjects “turaco” in the question while omitting crucial information about the objects (foods) in the documents, resulting the loss of essential semantic information. This motivates us to rethink the following question:
@@
StreamingLLM (Xiao et al., 2024)
H2O (Zhang et al., 2023)
SnapKV (Li et al., 2024)
PyramidInfer (Yang et al., 2024b)
PyramidKV (Cai et al., 2024)
@@
In light of this, we observe that the complete semantic information usually appear in a continuous sequence (Fang & Xie, 2022). Thus, we introduce a straightforward yet effective ChunkKV, grouping the tokens in a chunk as a basic compressing unit, which should be preserved or discarded as a whole. Thus, it retains the most informative semantic chunks from the original KV cache. As shown in Figure 1, preserving a chunk helps to catch the subject, predicate, and object. Furthermore, we investigate that the preserved KV cache indices by ChunkKV exhibit a higher similarity compared to previous methods. Consequently, we develop a technique called layer-wise index reuse, which reduces the additional computational time introduced by the KV cache compression method. As outlined in Table 1, recent highly relevant KV cache compression methods lack the ability to retain semantic information and efficiently reuse indices.
To evaluate ChunkKV’s performance, we conduct comprehensive experiments across multiple cutting-edge long-context benchmarks: long-context tasks including LongBench (Bai et al., 2024) and Needle-In-A-HayStack (NIAH) (Kamradt, 2023), in-context learning tasks such as GSM8K (Cobbe et al., 2021) and JailbreakV (Luo et al., 2024). And also different models including DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025),LLaMA-3-8B-Instruct (Meta, 2024), Mistral-7B-Instruct (Jiang et al., 2023a), and Qwen2-7B-Instruct (Yang et al., 2024a). Our experimental results demonstrate that ChunkKV surpasses existing KV cache compression methods in both efficiency and accuracy, primarily due to its ability to preserve essential information through selective chunk retention. These findings establish ChunkKV as a simple yet effective approach to KV cache compression.
@@
  • We propose ChunkKV, a simple KV cache compression method that uses the fragmentation method that keeps the semantic information, and propose the layer-wise index reuse technique to reduce the additional computational time.
@@ 2 Related Work
KV Cache Compression. KV cache compression technology has developed rapidly in the era of LLM, with methods mainly focused on evicting unimportant tokens. The compression process occurs before the attention blocks, optimizing both the prefilling time and GPU memory. Xiao et al. (2024) and Han et al. (2024) propose that initial and recent tokens consistently have high attention scores between different layers and attention heads. As a result, retaining these tokens in the KV cache is more likely to preserve important information. Furthermore, FastGen (Ge et al., 2023) evicts tokens based on observed patterns. H2O (Zhang et al., 2023) and SnapKV (Li et al., 2024) employ dynamic KV cache compression methods, evaluating the importance of tokens based on attention scores and then evicting the less important ones. As inference scenarios become increasingly complex, dynamic KV cache compression methods demonstrate powerful performance. Recently, Yang et al. (2024b) and Cai et al. (2024) have closely examined the distributions of attention scores during the pre-filling stage of the Retrieval-Augmented Generation (RAG) task, discovering a pyramidal KV cache compression pattern in different transformer layers.
@@
Chunking Method. The chunking methodology is widely used in the field of NLP due to its simplicity and effectiveness (Tjong Kim Sang & Veenstra, 1999). In the era of LLMs, chunking is primarily applied in data pre-processing. For example, Shi et al. suggest grouping related training data into chunks to achieve better convergence curves to pre-train LLMs. Fei et al. (2024) apply a topic-based chunking method to improve the semantic compression of prompts. Furthermore, chunking plays an important role in the Retrieval-Augmented Generation (RAG) field (Yepes et al., 2024; Smith & Troynikov, 2024; Anthropic, 2024). It serves to divide documents into units of information with semantic content suitable for embedding-based retrieval and processing by LLMs.
Layer-Wise Technique The layer-wise technique is widely used in the training and inference of large language models (LLMs). LISA (Pan et al., 2024a) is a layer-wise sampling method based on observations of the training dynamics of Low-Rank Adaptation (LoRA)(Hu et al., 2022) across layers. LAMB(You et al., 2020) is a layer-wise adaptive learning rate method that speeds up LLM training by stabilizing training convergence with large batch sizes. DoLa (Chuang et al., 2023) employs layer-wise contrasting to reduce hallucinations during LLM inference.
@@ 3.1 Preliminary Study of KV Cache Compression
With the increasing long-context capabilities of LLMs, the KV cache has become crucial for improving inference efficiency. However, it can consume significant GPU memory when handling long input contexts. The GPU memory cost of the KV cache for the decoding stage can be calculated as follows:
@@
where $B$ is the batch size, $S$ is the sequence length of prompt and decoded length, $L$ is the number of layers, $N$ is the number of attention heads, $D$ is the dimension of each attention head, and the first $2$ accounts for the KV matrices, while the last $2$ accounts for the precision when using float16. Table E shows the configuration parameters for LLaMA-3-8B-Instruct (Meta, 2024). With a batch size $B=1$ and a sequence length of prompt $S=2048$, the GPU memory cost of the KV cache is nearly $1$ GB. If the batch size exceeds 24, the GPU memory cost of the KV cache will exceed the capacity of an RTX 4090 GPU. To address this issue, KV cache compression methods have been proposed, with the aim of retaining only a minimal amount of KV cache while preserving as much information as possible. For more details on the LLM configuration parameters, refer to Appendix E.
@@ 3.2 Chunk Based KV Compression
To address the limitations of existing KV cache compression methods, we propose ChunkKV, a novel KV cache compression method that retains the most informative semantic chunks. The key idea behind ChunkKV is to group tokens in the KV cache into chunks that preserve more semantic information, such as a chunk containing a subject, verb and object. As illustrated in Figure 1, ChunkKV preserves the chunks of the KV cache that contain more semantic information. First, we define a chunk as a group of tokens that contain related semantic information. By retaining the most informative chunks from the original KV cache, ChunkKV can effectively reduce the memory usage of the KV cache while preserving essential information.
@@
  $\mathbf{A}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}\mathbf{K}^{T}$Attention scores for the observe window}
  $C\leftarrow\left\lceil\frac{T_{k}}{c}\right\rceil$Calculate the number of chunks}
@@
     $\mathbf{A}_{i}\leftarrow\sum_{j=(i-1)c+1}^{ic}\mathbf{A}_{:,j}$Sum of observation scores for each chunk}
@@
The Algorithm 1 shows the pseudocode outline of ChunkKV. First, following H2O (Zhang et al., 2023) and SnapKV (Li et al., 2024), we set the observe window by computing the observation scores $\mathbf{A}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}\mathbf{K}^{T}$, where $\mathbf{Q}_{T_{q}-w:T_{q}}$ is the observe window, $\mathbf{K}$ is the Key matrix and the window size $w$ is usually set to $\{4,8,16,32\}$. Next, the number of chunks $C$ is calculated as $C=\left\lceil\frac{T_{k}}{c}\right\rceil$, where $T_{k}$ is the length of the Key matrix and $c$ is the chunk size. The observation scores for each chunk are then computed as $\mathbf{A}_{i}=\sum_{j=(i-1)c+1}^{ic}\mathbf{A}_{:,j}$ for $i=1,2,\ldots,C$. Referring to previous works (Zhang et al., 2023; Li et al., 2024; Yang et al., 2024b; Cai et al., 2024), we still use the top-$k$ algorithm as ChunkKV’s sampling policy. For the top-$k$ chunk selection, the top-$k$ chunks are selected based on their observation scores, where $k=\left\lfloor\frac{L_{\text{max}}}{c}\right\rfloor$, and $L_{\text{max}}$ is the maximum length of the compressed KV cache. The size of the last chunk will equal $\text{min}(c,L_{\text{max}}-(k-1)\times c)$. The indices of the top-$k$ chunks will keep the original sequence order. In the compression step, the key and value matrices are only retained based on the selected indices, resulting in the compressed KV cache. Finally, the observe window of the original KV cache will be concatenated to the compressed KV cache by replacing the last $w$ tokens to keep important information. The compressed KV cache is then used for subsequent attention computations.
@@
Figure 2: Layer-wise similarity heatmaps of the preserved KV cache indices by SnapKV (left) and ChunkKV (right) on LLaMA-3-8B-Instruct.
3.3 Layer-Wise Index Reuse
Furthermore, we investigated the preserved KV cache indices by ChunkKV and found that they exhibit higher similarity compared to previous methods. Figure 2 shows the layer-wise similarity heatmaps of SnapKV and ChunkKV. Each cell represents the similarity between the preserved KV cache indices of two layers, with deeper colors indicating higher similarity. The results demonstrate that the KV cache indices preserved by ChunkKV are more similar to those in neighboring layers. As shown in Table 2, ChunkKV consistently achieves a higher average Jaccard similarity between adjacent layers compared to SnapKV in different model architectures, indicating that the retained token index in ChunkKV is more similar to each other. For a more detailed visualization, please refer to Appendix B.1.2.
Method H2O SnapKV ChunkKV
LLaMA-3-8B 25.31% 27.95% 57.74%
Qwen2-7B 14.91% 16.50% 44.26%
Mistral-7B 15.15% 15.78% 52.16%
Based on the above findings, we propose a training-free layer-wise index reuse method to further reduce the additional cost of the KV cache compression time, which reuses compressed token indices across multiple layers. This procedure is formally described in Algorithm 2. The ChunkKV compression process returns the compressed KV cache and their respective token indices, denoted as $\mathcal{I}_{l}$. For layer-wise index reuse, we define a grouping of layers such that all $N_{\text{reuse}}$ layers share the same token indices for ChunkKV. Specifically, for a group of layers $\left\{l,l+1,\ldots,l+N_{\text{reuse}}-1\right\}$, we perform ChunkKV on the first layer $l$ to obtain the token indices $\mathcal{I}_{l}$ and reuse $\mathcal{I}_{l}$ for the subsequent layers $l+1,l+2,\ldots,l+N_{\text{reuse}}-1$. The notation $\mathbf{K}_{l}[\mathcal{I}_{l}]$ and $\mathbf{V}_{l}[\mathcal{I}_{l}]$ indicates the selection of key and value caches based on the indices in $\mathcal{I}_{l}$. The efficiency analysis for layer-wise index reuse is provided in Appendix B.1.1.
Theoretical Understanding. We provide a theoretical understanding from the in-context learning (ICL) (Fang & Xie, 2022) to interpret why maintaining KV cache according to a continuous sequence in ChunkKV is better than according to sparse tokens. Informally speaking, the continuously chunk-level KV cache preserves the whole examples (semantic information) in ICL, thus reducing the requirement on distinguishability, i.e lower bound of KL divergence between the example and the question (Equation 4 in Condition 2). The complete analysis is provided in Appendix C.
4 Experiment Results
In this section, we conduct experiments to evaluate the effectiveness of ChunkKV on KV cache compression in two benchmark fields, with a chunk size set to 10 even for various model architectures. The first is the In-Context Learning benchmark, for which we select GSM8K (Cobbe et al., 2021) and Jailbreakv (Luo et al., 2024) to evaluate the performance of ChunkKV, furthermore we also include multi-step reasoning LLM DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025) to evaluate the performance of ChunkKV. The In-Context Learning scenario is a crucial capability for LLMs and has been adapted in many powerful technologies such as Chain-of-Thought (Wei et al., 2022; Diao et al., 2024; Pan et al., 2024b). The second is the Long-Context benchmark, which includes LongBench (Bai et al., 2024) and Needle-In-A-HayStack (NIAH) (Kamradt, 2023), both widely used for assessing KV cache compression methods. All experiments were conducted three times, using the mean score to ensure robustness.
4.1 In-Context Learning
The In-Context Learning (ICL) ability significantly enhances the impact of prompts on large language models (LLMs). For example, the Chain-of-Thought approach (Wei et al., 2022) increases the accuracy of the GSM8K of the PaLM model (Chowdhery et al., 2022) from 18% to 57% without additional training. In this section, we evaluate the performance of ChunkKV on the GSM8K, Many-Shot GSM8K (Agarwal et al., 2024), and JailbreakV (Luo et al., 2024) benchmarks.
@@ 4.1 In-Context Learning
Ratio StreamingLLM H2O SnapKV PyramidKV ChunkKV (Ours)
@@
GSM8K In the in-context learning scenario, we evaluated multiple KV cache compression methods for GSM8K (Cobbe et al., 2021), which contains more than 1,000 arithmetic questions on LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct (Meta, 2024), Qwen2-7B-Instruct (Yang et al., 2024a) and DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025). Follow the Agarwal et al. (2024), we consider many-shot GSM8K as a long-context reasoning scenario, which is a more challenging task than LongBench (Bai et al., 2024). The CoT prompt settings for this experiment are the same as those used by Wei et al. (2022), for many-shot GSM8K we set the number of shots to 50, which the prompt length is more than 4k tokens. For more details on the prompt settings, please refer to the APPENDIX G.
Table 4: Many-Shot GSM8K Performance Comparison.
Ratio StreamingLLM H2O SnapKV PyramidKV ChunkKV (Ours)

### v3

@@ Abstract
Large Language Models (LLMs) require significant GPU memory when processing long texts, with the key value (KV) cache consuming up to 70% of total memory during inference. Although existing compression methods reduce memory by evaluating the importance of individual tokens, they overlook critical semantic relationships between tokens, resulting in fragmented context and degraded performance. We introduce ChunkKV, which fundamentally reimagines KV cache compression by treating semantic chunks - rather than isolated tokens - as basic compression units. This approach preserves complete linguistic structures and contextual integrity, ensuring that essential meaning is retained even under aggressive compression. Our innovation includes a novel layer-wise index reuse technique that exploits the higher cross-layer similarity of preserved indices in ChunkKV, reducing computational overhead and improving throughput by 26.5%. Comprehensive evaluations on challenging benchmarks: LongBench, Needle-In-A-HayStack, GSM8K, and JailbreakV demonstrate that ChunkKV outperforms state-of-the-art methods by up to 8.7% in precision while maintaining the same compression ratio. These results confirm that semantic-aware compression significantly enhances both efficiency and performance for long-context LLM inference, providing a simple yet effective solution to the memory bottleneck problem.
**footnotetext: Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Corresponding Author.
@@ 1 Introduction
Large Language Models (LLMs) have become essential for addressing various downstream tasks of natural language processing (NLP), including summarization and question answering, which require the interpretation of a long context from sources such as books, reports, and documents, often encompassing tens of thousands of tokens [1, 2, 3]. Recent advances in long-context technology within the field of machine learning (ML) systems [4, 5, 6] have significantly improved computational throughputs and reduced latency of LLMs to process increasingly large input context lengths [7, 8] with historical KV cache (key value attentions). However, the memory requirement of the KV cache in serving super-long contexts becomes a new bottleneck [9, 10]. For instance, the KV cache for a single token in a 7B-parameter model requires approximately 0.5 MB of GPU memory, resulting in a 10,000-token prompt consuming around 5 GB of GPU memory.
@@
To address the substantial GPU memory consumption caused by KV caching, recent studies consider compressing the KV cache by pruning non-important discrete parts from the prompt tokens [9, 11, 12, 13, 14, 15, 16, 17]. H2O [9] and SnapKV [11] have shown that retaining less than 50% of the discrete KV cache can significantly reduce GPU memory usage with minimal impact on performance. However, we identify that the previous KV cache compression methods [9, 13] measure token importance isolatedly, neglecting the dependency between different tokens on the characteristics of the real-world language. For example, as shown in Figure 1, focusing on the importance of the token level can excessively focus on words about subjects “turaco” in the question while omitting crucial information about objects (foods) in the documents, resulting in the loss of essential semantic information. This motivates us to rethink the following question:
@@
StreamingLLM [6]
H2O [9]
SnapKV [11]
PyramidInfer [15]
PyramidKV [13]
@@
In light of this, we observe that complete semantic information usually appears in a continuous sequence [18, 19, 20]. Thus, we introduce a straightforward yet effective ChunkKV, grouping the tokens in a chunk as a basic compressing unit, which should be preserved or discarded as a whole. Thus, it retains the most informative semantic chunks from the original KV cache. As shown in Figure 1, preserving a chunk helps catch the subject, predicate, and object. In Table 12, we quantify the minimal loss and higher recovery of the attention score achieved by the chunk-based approach at the inference stage. Furthermore, we investigate that the preserved KV cache indices by ChunkKV exhibit a higher similarity compared to previous methods. Consequently, we develop a technique called layer-wise index reuse, which reduces the additional computational time introduced by the KV cache compression method. As outlined in Table 1, recent highly relevant KV cache compression methods lack the ability to retain semantic information and efficiently reuse indices.
To evaluate ChunkKV’s performance, we conduct comprehensive experiments across multiple cutting-edge long-context benchmarks: long-context tasks including LongBench [21] and Needle-In-A-HayStack (NIAH) [22], in-context learning tasks such as GSM8K [23] and JailbreakV [24]. And also different models including DeepSeek-R1-Distill-Llama-8B [25],LLaMA-3-8B-Instruct [26], Mistral-7B-Instruct [27], and Qwen2-7B-Instruct [28]. Our experimental results demonstrate that ChunkKV surpasses existing KV cache compression methods in both efficiency and accuracy, primarily due to its ability to preserve essential information through selective chunk retention. These findings establish ChunkKV as a simple yet effective approach to KV cache compression.
@@
  • We propose ChunkKV, a simple KV cache compression method that uses the fragmentation method that retains semantic information, and design the layer-wise index reuse technique to reduce the additional computational time.
@@ 2 Related Work
KV Cache Compression. KV cache compression technology has developed rapidly in the era of LLM, with methods mainly focused on evicting unimportant tokens. The compression process occurs before the attention blocks, optimizing both the prefilling time and GPU memory. Xiao et al. [6] and Han et al. [29] propose that initial and recent tokens consistently have high attention scores between different layers and attention heads. As a result, retaining these tokens in the KV cache is more likely to preserve important information. Furthermore, FastGen [12] evicts tokens based on observed patterns. H2O [9] and SnapKV [11] employ dynamic KV cache compression methods, evaluating the importance of tokens based on attention scores and then evicting the less important ones. As inference scenarios become increasingly complex, dynamic KV cache compression methods demonstrate powerful performance. Recently, Yang et al. [15] and Cai et al. [13] have closely examined the distributions of attention scores during the pre-filling stage of the Retrieval-Augmented Generation (RAG) task, discovering a pyramidal KV cache compression pattern in different transformer layers.
@@ 3.1 Preliminary Study of KV Cache Compression
With the increasing long-context capabilities of LLMs, the KV cache has become crucial for improving the inference efficiency. However, it can consume significant GPU memory when handling long input contexts. The GPU memory cost of the KV cache for the decoding stage can be calculated as follows:
@@
where $B$ is the batch size, $S$ is the sequence length of prompt and decoded length, $L$ is the number of layers, $N$ is the number of attention heads, $D$ is the dimension of each attention head, and the first $2$ accounts for the KV matrices, while the last $2$ accounts for the precision when using float16. Table E shows the configuration parameters for LLaMA-3-8B-Instruct [26]. With a batch size $B=1$ and a sequence length of prompt $S=2048$, the GPU memory cost of the KV cache is nearly $1$ GB. If the batch size exceeds 24, the GPU memory cost of the KV cache will exceed the capacity of an RTX 4090 GPU. To address this issue, KV cache compression methods have been proposed, with the aim of retaining only a minimal amount of KV cache while preserving as much information as possible. For more details on the LLM configuration parameters, refer to Appendix E.
@@ 3.2 Chunk Based KV Compression
To address the limitations of existing KV cache compression methods, we propose ChunkKV, a novel KV cache compression method that retains the most informative semantic chunks. The key idea behind ChunkKV is to group tokens in the KV cache into chunks that preserve more semantic information, such as a chunk containing a subject, verb, and object. As illustrated in Figure 1, ChunkKV preserves the chunks of the KV cache that contain more semantic information. First, we define a chunk as a group of tokens that contain related semantic information. By retaining the most informative chunks from the original KV cache, ChunkKV can effectively reduce the memory usage of the KV cache while preserving essential information. For more information on ChunkKV, please refer to Section A.
@@
$\mathbf{A}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}\mathbf{K}^{T}$triangleright Attention scores for the observe window
$C\leftarrow\left\lceil\frac{T_{k}}{c}\right\rceil$triangleright Calculate the number of chunks
@@
     $\mathbf{A}_{i}\leftarrow\sum_{j=(i-1)c+1}^{ic}\mathbf{A}_{:,j}$triangleright Sum of attention scores for each chunk
@@
The algorithm 1 shows the pseudocode for ChunkKV. First, following the approach of H2O [9] and SnapKV [11], we set the observe window by computing the attention scores $\mathbf{A}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}\mathbf{K}^{T}$, where $\mathbf{Q}_{T_{q}-w:T_{q}}$ is the observe window, $\mathbf{K}$ is the Key matrix and the window size $w$ is usually set to $\{4,8,16,32\}$. Next, the number of chunks $C$ is calculated as $C=\left\lceil\frac{T_{k}}{c}\right\rceil$, where $T_{k}$ is the length of the Key matrix and $c$ is the chunk size. The attention scores for each chunk are then computed as $\mathbf{A}_{i}=\sum_{j=(i-1)c+1}^{ic}\mathbf{A}_{:,j}$ for $i=1,2,\ldots,C$. We use the top-$k$ algorithm as the sampling policy for ChunkKV. The top-$k$ chunks are selected based on their attention scores, where $k=\left\lfloor\frac{L_{\text{max}}}{c}\right\rfloor$, and $L_{\text{max}}$ is the maximum length of the compressed KV cache. The size of the last chunk is set to $\text{min}(c,L_{\text{max}}-(k-1)\times c)$. The indices of the top-$k$ chunks preserve the original sequence order. In the compression step, only the key and value matrices corresponding to the selected indices are retained, resulting in the compressed KV cache. Finally, the observe window of the original KV cache will be concatenated to the compressed KV cache by replacing the last $w$ tokens to keep important information. The compressed KV cache is then used for subsequent attention computations. For implementation, we add vectorized operations, memory optimizations, etc., to optimize the code. Refer to Appendix A.5 for more details.
Figure 2: Layer-wise similarity heatmaps of the preserved KV cache indices by SnapKV (left) and ChunkKV (right) on LLaMA-3-8B-Instruct. Deep colors indicate higher similarity. More visualization can be found in Appendix B.1.3.
3.3 Layer-Wise Index Reuse
Method H2O SnapKV ChunkKV
LLaMA-3-8B 25.31% 27.95% 57.74%
Qwen2-7B 14.91% 16.50% 44.26%
Mistral-7B 15.15% 15.78% 52.16%
Furthermore, we investigated the KV cache indices preserved by ChunkKV and found that they exhibit higher similarity compared to previous methods. Figure 2 shows the layer-wise similarity heatmaps of SnapKV and ChunkKV. Each cell represents the similarity between the preserved KV cache indices of two layers, with deeper colors indicating higher similarity. The results demonstrate that the KV cache indices preserved by ChunkKV are more similar to those in neighboring layers.
@@ 3.3 Layer-Wise Index Reuse
As shown in Table 2, ChunkKV consistently achieves a higher average Jaccard similarity between adjacent layers compared to SnapKV in different model architectures, indicating that the retained token index in ChunkKV is more similar to each other. For a more detailed visualization, please refer to Appendix B.1.3.
Based on the above findings of the KV cache index, we propose a training-free layer-wise index reuse method to further reduce the additional cost of KV cache compression, which reuses compressed token indices across multiple layers. We evaluate the efficiency and effectiveness of the layer-wise index reuse method in Section 4.3. Reuse of the layer-wise index reduces the KV cache compression time by 20% compared to the FullKV baseline, with a performance drop of only 0.5%.
@@
# SLM=StreamingLLM, SKV=SnapKV, PKV=PyramidKV
Ratio SLM H2O SKV PKV ChunkKV (Ours)
@@
This layer-wise index reuse method is formally described in Algorithm 2. The ChunkKV compression process returns the compressed KV cache and their respective token indices, denoted as $\mathcal{I}_{l}$. For layer-wise index reuse, we define a grouping of layers such that all $N_{\text{reuse}}$ layers share the same token indices for ChunkKV. Specifically, for a group of layers $\left\{l,l+1,\ldots,l+N_{\text{reuse}}-1\right\}$, we perform ChunkKV on the first layer $l$ to obtain the token indices $\mathcal{I}_{l}$ and reuse $\mathcal{I}_{l}$ for the subsequent layers $l+1,l+2,\ldots,l+N_{\text{reuse}}-1$. The notation $\mathbf{K}_{l}[\mathcal{I}_{l}]$ and $\mathbf{V}_{l}[\mathcal{I}_{l}]$ indicates the selection of key and value caches based on the indices in $\mathcal{I}_{l}$. The efficiency analysis for layer-wise index reuse is provided in Appendix B.1.1.
Theoretical Understanding. We provide a theoretical understanding from the in-context learning (ICL) [20] to interpret why maintaining the KV cache according to a continuous sequence in ChunkKV is better than according to sparse tokens. Informally speaking, the continuously chunk-level KV cache preserves the whole examples (semantic information) in ICL, thus reducing the requirement on distinguishability, i.e., lower bound of KL divergence between the example and the question (Equation 4 in condition 2). The complete analysis is provided in Appendix C.
4 Experiment Results
In this section, we conduct experiments to evaluate the effectiveness of ChunkKV on KV cache compression in two benchmark fields, with a chunk size set to 10 even for various model architectures. The first is the In-Context Learning benchmark, for which we select GSM8K [23] and Jailbreakv [24] to evaluate the performance of ChunkKV, furthermore, we also include multi-step reasoning LLM DeepSeek-R1-Distill-Llama-8B [25] to evaluate the performance of ChunkKV. The In-Context Learning scenario is a crucial capability for LLMs and has been adapted in many powerful technologies such as Chain-of-Thought [30, 31, 32]. The second is the Long-Context benchmark, which includes LongBench [21] and Needle-In-A-HayStack (NIAH) [22], both widely used for assessing KV cache compression methods. All experiments were carried out three times, using the mean score to ensure robustness.
4.1 In-Context Learning
The In-Context Learning (ICL) ability significantly enhances the impact of prompts on LLMs. For example, the Chain-of-Thought approach [30] increases the accuracy of the GSM8K of the PaLM model [33] from 18% to 57% without additional training. In this section, we evaluate the performance of ChunkKV on the GSM8K, Many-Shot GSM8K [34], and JailbreakV [24] benchmarks.
Table 4: Many-Shot (50-shot) GSM8K Performance Comparison.
Ratio SLM H2O SKV PKV ChunkKV (Ours)